OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of –2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x). I dont get it, can anyone please help me?

OpenStudy (amistre64):

(-2+2) (3-3)^2 (x+2) (x-3)^2

OpenStudy (anonymous):

can you please explain?

OpenStudy (amistre64):

which part? what a zero is? how to form a zero from roots? why zeroed roots equal zero? or some other ?

OpenStudy (amistre64):

if i knew where your confusion lie, I could try to help you thru it

OpenStudy (asnaseer):

if a polynomial has three roots, lets call them \(r_1, r_2, r_3\), then it can be expressed as:\[f(x)=(x-r_1)(x-r_2)(x-r_3)=0\]

OpenStudy (amistre64):

they might be back ....

OpenStudy (asnaseer):

in your case, two of the roots are the same, so lets sat \(r_3=r_2\), this would lead to:\[f(x)=(x-r_1)(x-r_2)^2=0\] which is what amistre was showing you.

OpenStudy (amistre64):

the latex freezez up the IE

OpenStudy (asnaseer):

and on chrome as well - I get round it for now by entering a newline after I have typed it up - that seems to force it to render.